OpenStudy (anonymous):

Simplify the following expression. sqrt -49 sqrt -12

OpenStudy (callisto):

Is your question like this? \[\sqrt{-49}\sqrt{-12}\]

OpenStudy (anonymous):

yes please show me the steps to solve

OpenStudy (callisto):

If your question is that one... then \[\sqrt{-49}\sqrt{-12}\]\[=\sqrt{49 \times -1}\sqrt{12\times -1}\]\[=7\sqrt{-1}\times 2\sqrt{3\times -1}\]\[ = 7i \times 2\sqrt3i\]\[ = 7 \times 2\sqrt3 \times i^2\]\[=14\sqrt3 (-1)\]\[=-14\sqrt3\]

OpenStudy (anonymous):

I messed that one up. only got up to the 2nd line.. thank you

OpenStudy (anonymous):

I will go back and review why I missed those steps. appreciate it:)

OpenStudy (callisto):

Welcome :)